OpenStudy (anonymous):

Determine whether parallelogram JKLM with vertices J(-3, -2), K(2, -2), L(5, 2) and M(0, 2) is a rhombus, square, rectangle or all three.

OpenStudy (anonymous):

see all squares are rectangles ..

OpenStudy (anonymous):

first find the length of all the sides say JK=sqrt((-3-2)^2+(-2+2)^2)= sqrt(25)=5 units

OpenStudy (anonymous):

properties of a rhombus are same as a parallelogram bisecting angles must be perpendicular

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

the sides of a rhombus also have to have to same measurement also

OpenStudy (anonymous):

check the slope of JK and KL slope of JK= 0 ( parallel to xaxis) slope of kl =(2+2)/(5-2)=4/3 (not parallel to y axis ) i.e JK and KL ar not at right angles see that all the sides are of same length and the interior angles are not right angle hence only rhombus

OpenStudy (anonymous):

u can optionally find the length of diagonals which may be of different length in the case of rhombus but same in the case of squares and rectangles...